# Listening to the Noise: Blind Denoising with Gibbs Diffusion

### v1

@@ 1 Introduction
Denoising problems can be seen as a subset of the broader category of linear inverse problems, for which observations read ${\bm{y}}={\bm{A}}{\bm{x}}+{\bm{\varepsilon}}$ with ${\bm{A}}$ a general linear operator. Diffusion models have proven to be of great interest for these problems, generating a wealth of literature adjacent to our problem. In various techniques, diffusion-based priors have been integrated as plug-and-play models (see e.g., Kadkhodaie & Simoncelli, 2021; Meng & Kabashima, 2022; Kawar et al., 2022; Chung et al., 2023; Song et al., 2023a; Zhu et al., 2023; Rout et al., 2023), relying on the decomposition of the conditional score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t}\,|\,{\bm{y}})$. These methods typically involve approximations of the guiding term $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}\,|\,{\bm{x}}_{t})$. Still employing the diffusion model as a plug-and-play prior but avoiding such approximations, Cardoso et al. (2024) introduced a sequential Monte Carlo method for Bayesian inverse problems. Interestingly, Murata et al. (2023) developed a Gibbs sampler to address the case of a linear operator ${\bm{A}}$ depending on unknown parameters ${\bm{\phi}}$. While the methodology bears similarities with the approach taken in this paper, it does not seem easily adaptable to the blind denoising problem we address in this work.
@@ 2.1 Diffusion Models as Denoising Posterior Samplers
[…]
@@ 2.2 Blind Denoising with Gibbs Diffusion
    Sampling of $p({\bm{x}}\,|\,{\bm{y}},{\bm{\phi}})$. Sect. 2.1 has shown that a diffusion model can be trained to naturally address posterior sampling of $p({\bm{x}}\,|\,{\bm{y}},{\bm{\phi}})$ provided that the forward process meets certain conditions. In practice, our forward process must take the form of Eq. (2) and $f$ and $g$ have to be such that $\sigma\leq b(1)/a(1)$, which is always easily achievable (cf Sect. 3). Moreover, since ${\bm{\phi}}$ is a priori unknown, we have to train this diffusion model conditionally to the noise parameters ${\bm{\phi}}$.
@@ 2.3 Properties of the Gibbs Sampler
In general, there may not exist a joint distribution over ${\bm{x}}$ and $\phi$, which matches the conditionals $p({\bm{\phi}}\mid{\bm{y}},{\bm{x}})$ and $q({\bm{x}}\mid{\bm{y}},{\bm{\phi}})$, in which case we say the conditionals are incompatible. Incompatibility can notably arise if $q({\bm{x}}\mid{\bm{y}},\phi)$ poorly approximates the true conditional $p({\bm{x}}\mid{\bm{y}},\phi)$ (e.g Hobert & Casella, 1998; Liu et al., 2012). We show in App. D that the main condition for compatibility is,
@@
We adapt this result from Arnold et al. (2001, Theorem 1) and emphasize in our formulation the relationship between the conditionals. Conceptually, Eq. (7) tells us that there can be no measurable set in which ${\bm{x}}$ is likely given $\phi$ but $\phi$ is unlikely given ${\bm{x}}$. If the compatibility condition is verified, then the stationary distribution is unique (Gourieroux & Montfort, 1979).
@@
Assume now compatibility and denote $\pi({\bm{\phi}},{\bm{x}}\mid{\bm{y}})$ the stationary distribution. How do the approximations in our Gibbs sampler impact our inference on ${\bm{\phi}}$? Let $\rho_{M}({\bm{\phi}}\mid{\bm{y}})$ be the distribution of ${\bm{\phi}}$ after $M$ sampling iterations and let $D_{\text{TV}}$ denote the total variation distance. Then,
@@
Suppose $p({\bm{\phi}}\mid{\bm{y}},{\bm{x}})$ and $q({\bm{x}}\mid{\bm{y}},{\bm{\phi}})$ are compatible, and are the conditionals of the joint $\pi({\bm{\phi}},{\bm{x}}\mid{\bm{y}})$. Then the Kullback-Leibler divergence is
[…]
@@
Without detailed knowledge of $q({\bm{x}}\mid{\bm{y}},\phi)$ and $p({\bm{x}},\phi\mid{\bm{y}})$, we cannot say more about the error in our Gibbs sampler using theory alone. Compatibility between conditionals has been studied in restricted contexts, for instance when the conditionals belong to certain parametric families (Arnold et al., 2001; Liu et al., 2012), but this does not describe our setting where the conditionals are intractable. Ergodicity and convergence rates for the Gibbs and HMC samplers are discussed in references (Geman & Geman, 1984; Livingstone et al., 2019) but cannot be computed explicitly.
@@
In practice, we empirically check convergence and the algorithm’s calibration. Convergence diagnostics help us detect if the Gibbs sampler has been run for a sufficient number of iterations. If poor calibration persists after convergence has been detected, then by Theorem 2.3, the diffusion model introduces a non-negligible error in our inference.
@@ 3.1 Denoising for Natural Images
ImageNet 0.06 31.0±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 30.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 29.3±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 32.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 33.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 33.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 34.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 34.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 33.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 32.3±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 35.3±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$
0.1 27.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 26.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.1}$ 26.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 29.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 31.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 29.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 32.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 32.1±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 31.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 29.9±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 32.9±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$
0.2 23.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 21.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.1}$ 23.0±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 25.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 28.1±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 28.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 26.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 29.3±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 29.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 28.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 27.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 30.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$
CBSD68 0.06 31.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 30.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.1}$ 29.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 32.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 33.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 34.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 34.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 35.0±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 34.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 32.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 35.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$
0.1 27.9±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 26.9±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.1}$ 26.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 29.1±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.3±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 28.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 31.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 33.0±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 32.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 30.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$ 33.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.4}$
0.2 23.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 21.7±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.1}$ 23.0±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 25.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 27.8±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 28.2±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 25.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 28.5±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 29.6±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ 28.9±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.2}$ 27.4±${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$ ${}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\scriptscriptstyle\pm 0.3}$
@@
Table 1: Denoising performance in terms of PSNR ($\uparrow$, in dB) for GDiff (blind) and baselines BM3D (non-blind) and DnCNN (blind). We report mean PSNR and standard error computed on batches of 50 images. For GDiff, we provide performance for both posterior samples ${\bm{x}}$ and estimates of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$.

### v2

@@ 1 Introduction
Denoising problems can be seen as a subset of the broader category of linear inverse problems, for which observations read ${\bm{y}}={\bm{A}}{\bm{x}}+{\bm{\varepsilon}}$ with ${\bm{A}}$ a general linear operator. Diffusion models have proven to be of great interest for these problems, generating a wealth of literature adjacent to our problem. In various techniques, diffusion-based priors have been integrated as plug-and-play models (see e.g., Kadkhodaie & Simoncelli, 2021; Meng & Kabashima, 2022; Kawar et al., 2022; Chung et al., 2023; Song et al., 2023a; Zhu et al., 2023; Rout et al., 2023), relying on the decomposition of the conditional score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t}\,|\,{\bm{y}})$. These methods typically involve approximations of the guiding term $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}\,|\,{\bm{x}}_{t})$. In Appendix E, we discuss limitations of some of these methods for our problem. Still employing the diffusion model as a plug-and-play prior but avoiding such approximations, Cardoso et al. (2024) introduced a sequential Monte Carlo method for Bayesian inverse problems. Interestingly, Murata et al. (2023) developed a Gibbs sampler to address the case of a linear operator ${\bm{A}}$ depending on unknown parameters ${\bm{\phi}}$. While the methodology bears similarities with the approach taken in this paper, it does not seem easily adaptable to the blind denoising problem we address in this work. Finally, let us mention Laroche et al. (2024), which addresses blind deconvolution problems by estimating the maximum a posteriori using an Expectation-Minimization algorithm combined with a guidance-based sampling method.
@@ 2.1 Diffusion Models as Denoising Posterior Samplers
${\bm{z}}_{t}=a(t){\bm{x}}+b(t){\bm{\varepsilon^{\prime}}},\hfill\text{with}~{}\varepsilon^{\prime}\sim\mathcal{N}(0,{\bm{\Sigma}}_{\bm{\varphi}}),$ (3)
@@ 2.2 Blind Denoising with Gibbs Diffusion
    Sampling of $p({\bm{x}}\,|\,{\bm{y}},{\bm{\phi}})$. Sect. 2.1 has shown that a diffusion model can be trained to naturally address posterior sampling of $p({\bm{x}}\,|\,{\bm{y}},{\bm{\phi}})$ provided that the forward process meets certain conditions. In practice, our forward process must take the form of Eq. (2) and $f$ and $g$ have to be such that $\sigma\leq b(1)/a(1)$, which is always easily achievable (cf Sect. 3). Moreover, since ${\bm{\phi}}$ is a priori unknown, we train a conditional diffusion model where ${\bm{\phi}}$ is given an an input to the score network, using a range of values consistent with the prior distribution $p({\bm{\phi}})$.
@@ 2.3 Properties of the Gibbs Sampler
In general, there may not exist a joint distribution over ${\bm{x}}$ and $\phi$, which matches the conditionals $p({\bm{\phi}}\mid{\bm{y}},{\bm{x}})$ and $q({\bm{x}}\mid{\bm{y}},{\bm{\phi}})$, in which case we say the conditionals are incompatible. Incompatibility can notably arise if $q({\bm{x}}\mid{\bm{y}},\phi)$ poorly approximates the true conditional $p({\bm{x}}\mid{\bm{y}},\phi)$ (e.g., Hobert & Casella, 1998; Liu et al., 2012). We show in App. D that the main necessary condition for compatibility is,
@@
We adapt this result from Arnold et al. (2001, Theorem 1) and emphasize in our formulation the relationship between the conditionals. Conceptually, Eq. (7) tells us that there can be no measurable set in which ${\bm{x}}$ is likely given $\phi$ but $\phi$ is unlikely given ${\bm{x}}$. If the compatibility condition is verified, then under mild conditions the stationary distribution is unique and we say the two conditionals verify determinacy; see Gourieroux & Montfort (1979) and Liu et al. (2021) for a more detailed discussion.
@@
Assume now compatibility and determinacy. Denote $\pi({\bm{\phi}},{\bm{x}}\mid{\bm{y}})$ the unique stationary distribution. How do the approximations in our Gibbs sampler impact our inference on ${\bm{\phi}}$? Let $\rho_{M}({\bm{\phi}}\mid{\bm{y}})$ be the distribution of ${\bm{\phi}}$ after $M$ sampling iterations and let $D_{\text{TV}}$ denote the total variation distance. Then,
@@
Suppose $p({\bm{\phi}}\mid{\bm{y}},{\bm{x}})$ and $q({\bm{x}}\mid{\bm{y}},{\bm{\phi}})$ are compatible and are the conditionals of the joint $\pi({\bm{\phi}},{\bm{x}}\mid{\bm{y}})$. Then the Kullback-Leibler divergence is
$\text{KL}[p({\bm{\phi}}\mid{\bm{y}})||\pi({\bm{\phi}}\mid{\bm{y}})]=\mathbb{E}_{p({\bm{\phi}}\mid{\bm{y}})}\log\mathbb{E}_{p({\bm{x}}\mid{\bm{y}})}\frac{q({\bm{x}}\mid{\bm{y}},{\bm{\phi}})}{p({\bm{x}}\mid{\bm{y}},{\bm{\phi}})}.$
@@
Without detailed knowledge of $q({\bm{x}}\mid{\bm{y}},\phi)$ and $p({\bm{x}},\phi\mid{\bm{y}})$, we cannot say more about the error in our Gibbs sampler using theory alone. Compatibility between conditionals has been studied in restricted contexts, for instance when the conditionals belong to certain parametric families (Arnold et al., 2001; Liu et al., 2012), but this does not describe our setting where the conditionals are intractable. Motivated by the variational autoencoder, Liu et al. (2021) recently proposed a loss function (eq. 2 in reference) which, when minimized to 0, enforces compatibility; checking the viability of this strategy in our setting is left as future work. Ergodicity and convergence rates for the Gibbs and HMC samplers are discussed in references (Geman & Geman, 1984; Livingstone et al., 2019) but typically, these rates cannot be computed explicitly.
@@
In practice, we recommend empirically checking convergence and the algorithm’s calibration. Convergence diagnostics help us detect if the Gibbs sampler has been run for a sufficient number of iterations and whether a stationary distribution exists. If poor calibration persists after convergence has been detected, then by Theorem 2.3, the diffusion model introduces a non-negligible error in our inference.
@@ 3.1 Denoising for Natural Images
ImageNet 0.06 31.0±0.2 (30.2)±0.2 29.3±0.3 32.2±0.3 33.7±0.3 33.4±0.3 31.5±0.3 34.4±0.3 34.7±0.4 (33.8)±0.4 32.3±0.4 35.3±0.4
0.1 27.8±0.2 (26.8)±0.1 26.7±0.2 29.4±0.2 31.8±0.3 31.8±0.4 29.7±0.4 32.7±0.4 32.1±0.4 (31.5)±0.3 29.9±0.4 32.9±0.3
0.2 23.5±0.2 (21.7)±0.1 23.0±0.3 25.7±0.3 28.1±0.4 28.4±0.4 26.5±0.4 29.3±0.4 29.5±0.4 (28.6)±0.4 27.6±0.4 30.5±0.4
CBSD68 0.06 31.2±0.2 (30.6)±0.1 29.2±0.2 32.2±0.2 33.8±0.3 34.2±0.3 31.2±0.3 34.4±0.3 35.0±0.3 (34.8)±0.3 32.2±0.3 35.5±0.3
0.1 27.9±0.2 (26.9)±0.1 26.2±0.3 29.1±0.3 31.3±0.3 31.7±0.3 28.6±0.3 31.8±0.3 33.0±0.3 (32.7)±0.3 30.6±0.4 33.8±0.4
0.2 23.5±0.2 (21.7)±0.1 23.0±0.3 25.6±0.2 27.8±0.3 28.2±0.3 25.4±0.3 28.5±0.3 29.6±0.3 (28.9)±0.2 27.4±0.3 30.6±0.3
@@
Table 1: Denoising performance in terms of PSNR ($\uparrow$, in dB) for GDiff (blind) and baselines BM3D (non-blind) and DnCNN (blind). We report mean PSNR and standard error computed on batches of 50 images. For GDiff, we provide performance for both posterior samples ${\bm{x}}$ and estimates of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$. We point out that DnCNN was trained with white noises only, hence results obtained for $\varphi\neq 0$ could be sub-optimal.